OpenStudy (anonymous):

Simplify each expression leaving only positive exponents. (I'm still not sure how to do these) (-2c^-2)(ac^3)(-4a^-4c) (3x^-3 y^-2)^-3 (2x^-5 y^3)^2/(4x^-3 y)^0 3^-1 x^2 y^-4/2^-2x^5 y^3

OpenStudy (anonymous):

\[(-2)c^{-2}ac^{-3}(-4)a^{-4}c \] in general \[X^aX^b= X^{a+b}\] so \[c^{-2}c^{-3}c^{1} = c^{-2-3+1}=c^{-4}\] so you have \[+8 c^{-4} a^{1}a^{-4}\] can you finish?

OpenStudy (anonymous):

Do you have to combine a^1 with a^-4?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

And then I will get the answer for that one?

OpenStudy (anonymous):

Almost! remember it says only positive exponents. so if you get \[a^{-100}\] that would be wrong. Remember though \[X^{-a} = \frac{1}{X^a}\] so \[a^{-100} = \frac{1}{a^{100}}\] would be correct.

OpenStudy (anonymous):

oooh okay thank you

OpenStudy (anonymous):

\[X^aX^b=X^{a+b}\] \[\left(x^a\right)^b = X^{ab}\] \[x^{-a}=\frac{1}{X^a}\] \[X^a +X^b \neq X^{a+b}\]

OpenStudy (anonymous):

So would i write it like this?\[8c ^\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

\[\frac{8}{c^4a^3}\]

OpenStudy (anonymous):

oh okay thanks.

OpenStudy (anonymous):

\[ X^{1/a} = \sqrt[a]{X}\]

OpenStudy (anonymous):

????